What would the perimeter of a rectangle be with a width of of 7 feet and a length of 8 feet?
Answer by stanbon(75887)

What would the perimeter of a rectangle be with a width of of 7 feet and a length of 8 feet?
----
Perimeter = 2(L + W)
P = 2(8 + 7) = 30 feet
=======================
Cheers,
Stan H.